The Real Estate Group NY reports that the mean monthly rent for a one-bedroom apartment without a doorman in Manhattan is $2554. Assume the standard deviation is $453. A real estate firm samples 96 apartments. What is the probability that the sample mean rent is less than $2,619.65?
Answer #1

Solution :

Given that ,

mean = = $2554

standard deviation = = $453

n = 96

= 2554

=  / n = 453 / 96=46.2341

P( < $2619.65) = P[( - ) / < (2619.65-2554) / 46.2341]

= P(z <1.42 )

Using z table  

= 0.9222

probability=0.9222